OpenStudy (anonymous):

Help find all real solutions of the equation sqroot2x+1 )+1=x

OpenStudy (abb0t):

Square both sides and set equal to zero.

OpenStudy (anonymous):

\[\sqrt{2x+1} +1=x\]

OpenStudy (anonymous):

so x^2+2x+2=0

OpenStudy (abb0t):

Did you do: \[(\sqrt{2x+1} +1)^2 = x^2\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i got \[x^2+2x+2=0\]

OpenStudy (anonymous):

helllo @abb0t

OpenStudy (anonymous):

hello :(